OpenStudy (anonymous):

In 1990, the life expectancy of females was 78.8 yr. In 2001, it was 79.8 yr. Let F(t) = life expectancy and t = the number of years since 1990. a) Find a linear function that fits the data. b) Use the function of part (a) to estimate the life expectancy of females in 2010.

OpenStudy (anonymous):

a. F(t)=78.8 + 1/11t b.80.6

OpenStudy (anonymous):

In 1990, the life expectancy of females was 78.8 yr. In 2001, it was 79.8 yr. Let F(t) = life expectancy and t = the number of years since 1990 (79.8-78.8)/(2001-1990) 1/11 F(0) = 78.8 F(t) = t/11 + b b = 78.8 F(t) = t/11 + 78.8 F(20) = 20/11 + 78.8 =80.6181